OpenStudy (anonymous):

Divide: 6x^3-4x^2 / 2x^2+1

OpenStudy (anonymous):

Long division

OpenStudy (anonymous):

Try factorising.

OpenStudy (confusionist):

Combine like terms. 6x-4x-2x= 0 2-1-3+2=0 The answer is 0.

OpenStudy (shamil98):

er no ^

OpenStudy (anonymous):

It's division how? @Confusionist

OpenStudy (ranga):

3x - 2 -------------- 2x^2 + 0x + 1 | 6x^3 - 4x^2 | 6x^3 + 3x --------------------- -4x^2 + 3x -4x^2 - 2 --------------------- 3x - 2 6x^3-4x^2 / 2x^2+1 = 3x - 2 + (3x - 2) / (2x^2 + 1)